Given MUx = X, MUY = 4Y, Price of X is 3 Birr per unit and price of Y is also 3 Birr per unit. If
the income of the consumer is 1200 Birr, find the amounts of X and Y that the consumer chooses
to consume so as to maximize his utility.
1
Expert's answer
2021-07-14T04:32:02-0400

Solution

"\frac{MU_x}{p_x}=\frac{MU_y}{P_y}"


"\frac{x}{3}=\frac{4y}{3}"


Budget line = I= PxX+PyY


1200=3x+3y


"\frac{MU_x}{MU_y}=\frac{P_x}{P_y}=\frac{x}{4y}"


x=4y


1200=3(4y)+3y

1200=15y

y=80


1200=3x+240

960=3x

x=320


Utility is maximized when


y=80

x=320

MUx= 320

MUy= 320